OpenStudy (anonymous):

-4x-6y-72 -4x+8y=-88

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

-4x-6y-72 -4x+8y=-88 -4x-4x-6y+8y-76=-88 -8x+2y=-12